# Diff-BBO: Diffusion-Based Inverse Modeling for Black-Box Optimization

### v1

@@ 5 Experiments
To validate the efficacy of Diff-BBO, we conduct experiments on six real-world online black-box optimization tasks for both continuous and discrete optimization tasks. Ablation studies are performed to verify the effectiveness of the proposed acquisition function and the robustness of our model in relation to the batch size. More details of the experimental setups are provided in Appendix D.
@@ 5.1 Dataset
We restructured $5$ high-dimensional real-world tasks from Design-Bench to facilitate online black-box optimization. We test on $3$ continuous and $2$ discrete tasks. In D’Kitty and Ant Morphology, the goal is to optimize for the morphology of robots. In Superconductor, the aim is to optimize for finding a superconducting material with a high critical temperature. TFBind8 and TFBind10 are discrete tasks where the goal is to find a DNA sequence that has a maximum affinity to bind with a specified transcription factor. We also include a Molecular Discovery task to optimize a compound’s activity against a biological target with therapeutic value. Each optimization iteration is allocated $100$ queries to the oracle function (batch size $N=100$), with a total of $16$ iterations conducted. More details of the dataset are provided in Section D.1.
@@ 5.2 Baselines
We consider four existing groups of baselines, including Bayesian optimization (BO), Likelihood-free BO (LFBO) (Song et al., 2022), Evolutionary algorithms (Brindle, 1980; Real et al., 2019), and random sampling. For BO approaches, we include Gaussian Processes (GP) with Monte Carlo (MC)-based batch Expected Improvement (EI) and MC-based batch Upper Confidence Bound (UCB) (Wilson et al., 2017) as the acquisition functions. For LFBO, we use EI and Probability of Improvement (PI) as the acquisition functions.
@@ 5.3 Results
Figure 3: Comparison of Diff-BBO with baselines for online black-box optimization on DesignBench and Molecular Discovery task. We plot the mean and standard deviation across three random seeds. Diff-BBO demonstrates superior performance with very few queries to the oracle.
Figure 3 illustrates the performance across six datasets for all baselines and our proposed algorithm. Notably, Diff-BBO consistently outperforms other baselines in both discrete and continuous settings, with the sole exception of the TF-BIND-8 task. Specifically, in the Ant and Dkitty tasks, Diff-BBO demonstrates a significant lead over all baseline methods, starting from the very first iteration of the online optimization process. This remarkable performance can be attributed to Diff-BBO’s diffusion model-based inverse modeling approach, which effectively learns the data manifold in the design space from the initial dataset, even when the initial dataset lacks data with high objective function values. In contrast, the forward approach employed by BO and LFBO, which relies solely on optimizing the trained surrogate model, is more prone to converging on suboptimal solutions.
@@ 5.4 Ablation study
In this section, we conduct ablation studies to investigate the impact of our designed acquisition function, UaE. We compare Diff-BBO with the fixed condition approach. Instead of using UaE to dynamically determine which $y$ to condition on, the fixed condition approach always generates new samples conditioned on $w\cdot\phi_{k}$ (Line 9 of Algorithm 1) with a fixed weight $w$. As shown in Figure 5, Diff-BBO consistently outperforms the fixed condition approach. This demonstrates that our acquisition function is effective in identifying the optimal $y$ for conditioning.
Figure 4: Impact of acquisition function design for black-box optimization on the TFBIND10 task. Comparison of Diff-BBO with five fixed-condition approaches, each with different conditioning weights. Results averaged across three random runs.
Figure 5: Ablation study to evaluate the effect of batch size on the superconductor task. The mean and standard deviation across three random seeds are plotted. Diff-BBO shows robust performances across different batch size given the same total number of evaluations.
Furthermore, we evaluate the effect of batch size, aka the number of queries per iteration on Diff-BBO on the Superconductor task. As shown in Figure 5, we compare the objective function score over number of function evaluations. We can see the performance of our approach remains similar when the batch size becomes small, suggesting remarkable robustness across different batch sizes. Hence, Diff-BBO is a highly-scalable inverse modeling approach that can efficiently leverage parallelism to handle larger computational loads without compromising performance.
6 Conclusion, Limitation, and Future Work
@@
In this paper, we introduced Diff-BBO, a novel inverse modeling approach for black-box optimization that leverages the uncertainty of conditional diffusion models. Our method enhances the scalability in high-dimensional BBO problems and effectively quantifies uncertainty in inverse modeling. By utilizing the novel acquisition function UaE, Diff-BBO strategically proposes objective function values to improve sample efficiency in online settings. Our empirical evaluations on the Design-Bench benchmark and molecular design experiments demonstrate that Diff-BBO achieves state-of-the-art performance, establishing its potential as a robust tool for efficient and effective black-box optimization. Theoretically, we prove that using UaE leads to optimal optimization solutions.
We conclude by discussing the limitations and potential extensions of Diff-BBO: (i) limitations with respect to diffusion modeling: While we benefit from diffusion modeling to learn the manifold, we also inherit its drawbacks, such as slower sampling times and increased computational demands. (ii) acquisition function improvement: Our current implementation for the acquisition function, UaE, requires presetting the candidate sets. This necessitates additional hyperparameter tuning. (iii) BBO extensions: Diff-BBO can be extended to various Bayesian optimization settings, including multi-objective and multi-fidelity Bayesian optimization (Wu et al., 2023; Li et al., 2020).
@@ Appendix B Analysis of Sub-optimality for Black-box Function
At each iteration $k\in[K]$, let $y_{k}^{*}$ be the target function value that we condition on, and $p_{\theta}$ be the model learned by the conditional diffusion model. We define the performance metric for active BBO problem, which is the sub-optimal performance gap between the expected value of $\bm{x}\sim p_{\theta}(\cdot|y)$ and the target function value $y^{*}$. It is formally described as follows:
$\Delta(p_{\theta},y_{k}^{*})=\left|y_{k}^{*}-\max_{j\in[N]}f(\bm{x}_{j})\right|,~{}~{}\text{where}\bm{x}_{j}\sim p_{\theta}(\cdot|y_{k}^{*},\mathcal{D}),~{}~{}\forall j\in[N].$ (20)
For simplicity of analysis, we consider $N=1$, and let the generated sample at the $k$-th iteration be $\bm{x}_{k}$ in the remaining text. We remark that all proofs go through smoothly for general $N$ with more nuanced notations, and do not affect the conclusions being drawn. To proceed with the proofs in this section, we first state the formal assumptions for the black-box function $f(\cdot)$.
@@
For $y_{k}^{*}$ at iteration $k$, since we assume the existentence of $\theta^{*}\sim p(\theta\mid\mathcal{D})$, and $\bm{x}_{j}$ is generated from $p_{\theta}(\cdot|y_{k}^{*},\mathcal{D})$, where $\theta\sim p(\theta\mid\mathcal{D})$ is the parameter of the diffusion model at the $k$-th iteration, we can express ${\mathbb{E}}\left[{\Delta(p_{\theta},y_{k}^{*})}\right]$ as follows:
${\mathbb{E}}\left[{\Delta(p_{\theta},y_{k}^{*})}\right]$ $={\mathbb{E}}\left[{\mathbb{E}}\left[|f(\bm{x}^{*})-f(\bm{x}_{j})|\mid\mathcal{D}\right]\right]$
$=\left|{\mathbb{E}}\left[{\mathbb{E}}\left[f(\bm{x}^{*}\mid\mathcal{D})\right]\right]-{\mathbb{E}}\left[\left[f(\bm{x}_{j})\mid\mathcal{D}\right]\right]\right|,$
where the first equality is by tower rule, and the second inequality is due to the fact that $\bm{x}^{*}$ and $\bm{x}$ are independent. Since the black-box function $f$ is measurable, and by the nature of Algorithm 1, the generated sample $\bm{x}$, the proposed value $y_{k}^{*}$, the predictive distribution $p_{\theta}(\cdot|y_{k}^{*},\mathcal{D})$, the posterior distribution $p(\theta\mid\mathcal{D})$ are $\sigma(\mathcal{D})$-measurable at iteration $k$, the only randomness in $f(\bm{x})$ comes from the random sampling in the algorithm. Thus conditioned on the training data $\mathcal{D}$,
${\mathbb{E}}\left[f(\bm{x}_{j})\mid\mathcal{D}\right]$ $={\mathbb{E}}\left[f(\bm{x}_{j})\mid\mathcal{D},p(\theta,\mathcal{D})\right]$
$=\int_{\bm{x}}\int_{\theta}f(\bm{x})p_{\theta}(\cdot|y_{k}^{*},\mathcal{D})p(\theta|\mathcal{D})\mathop{}\!\mathrm{d}\theta\mathop{}\!\mathrm{d}\bm{x}$
$={\mathbb{E}}\left[f(\bm{x}^{*})\mid\mathcal{D}\right].$
Taking expectation of both sides completes the proof. ∎
@@
At every iteration $k\in[K]$, let the proposed function value that the conditional diffusion model conditions on be $y_{k}^{*}$. By Section 4.4, Conditioned on $y_{k}^{*}$, which is $\sigma(\mathcal{D})$-measurable, taking the variance of both sides of Equation 20 leads to,
$\mathrm{Var}_{\theta_{i}\sim p(\cdot\mid\mathcal{D})}\left(\Delta(p_{\theta},y_{k}^{*})\right)$ $={\mathbb{E}}\left[\left({\Delta(p_{\theta},y_{k}^{*})}\right)^{2}\right]-\left({\mathbb{E}}\left[{\Delta(p_{\theta},y_{k}^{*})}\right]\right)^{2}={\mathbb{E}}\left[\left({\Delta(p_{\theta},y_{k}^{*})}\right)^{2}\right].$
As a result, we only need to bound the second-order moment of $\Delta(p_{\theta},y_{k}^{*})$ with the $\Delta_{\mathrm{epidemic}}$. Under 2,
${\mathbb{E}}\left[\left({\Delta(p_{\theta},y_{k}^{*})}\right)^{2}\right]$ $={\mathbb{E}}\left[\left(y_{k}^{*}-f(\bm{x}_{j})\right)^{2}\right]$
$={\mathbb{E}}\left[\left(f(\bm{x}_{k}^{*})-f(\bm{x}_{k})\right)^{2}\right]$
$\leq L^{2}{\mathbb{E}}\left[\|\bm{x}_{k}^{*}-\bm{x}_{k}\|^{2}\right].$
Since the black-box function $f$ is measurable, and by the nature of Algorithm 1, the generated sample $\bm{x}_{k}$, the proposed value $y_{k}^{*}$, the predictive distribution $p_{\theta}(\cdot|y_{k}^{*},\mathcal{D})$, the posterior distribution $p(\theta\mid\mathcal{D})$ are $\sigma(\mathcal{D})$-measurable at iteration $k$. Thus conditioned on the training data $\mathcal{D}$, with a similar argument in the proof of Section 4.4, we have
${\mathbb{E}}[\bm{x}_{k}^{*}]={\mathbb{E}}[\bm{x}_{k}].$
As a result,
$\mathrm{Var}\left(\Delta(p_{\theta},y_{k}^{*})\right)\leq L^{2}\mathrm{Var}\left(\bm{x}_{k}^{*}-\bm{x}_{k}\right).$ (21)
Finally, by linearity of expectation,

### v3

@@ 6 Experiments
Figure 2: Comparison of Diffusion-BBO with baselines for online black-box optimization on DesignBench and Molecular Discovery task. All plots start at iteration 1 after one round of data queries. We plot the mean values and the confidence interval based on three random runs. Diffusion-BBO exhibits superior performance with respect to sample efficiency.
To validate the efficacy of Diffusion-BBO, we conduct experiments on six online black-box optimization tasks for both continuous and discrete optimization settings. We arrange the pre-collected data in ascending order based on objective values and select data from the 25th to the 50th percentile as the initial training dataset. The goal is to demonstrate that low-quality data, with below-average objective function values, is sufficient for the initial training stage of Diffusion-BBO. Each optimization iteration is allocated $100$ queries to the oracle function (batch size $N=100$), with a total of $16$ iterations conducted. For the fixed set of weights, we use $\mathcal{W}=\{0.6,0.7,0.8,0.9,1.0\}$. The corresponding candidate set then becomes $\mathcal{Y}_{k}=\{y_{k}=w\cdot\phi_{k}\mid w\in\mathcal{W}\}$. The rationale behind this choice is that $w>1$ falls outside the training data distribution, which we also verify through the ablation study. For UaE computation, we take the logarithm of both terms to ensure they are within the same range. More details of the experimental setup are provided in Appendix D.
@@ 6.1 Dataset
We restructured $5$ high-dimensional scientific discovery tasks from Design-Bench to facilitate online black-box optimization. We test on $3$ continuous and $2$ discrete tasks. In D’Kitty and Ant Morphology, the goal is to optimize for the morphology of robots. In Superconductor, the aim is to optimize for finding a superconducting material with a high critical temperature. TFBind8 and TFBind10 are discrete tasks where the goal is to find a DNA sequence that has a maximum affinity to bind with a specified transcription factor. We also include the Molecular Discovery task to optimize a compound’s activity against a biological target with therapeutic value. More details of the dataset are provided in Section D.1.
Figure 3: Impact of acquisition function design for black-box optimization on both discrete task (TFBind10) and continous task (D’Kitty). Comparison of Diffusion-BBO with UaE against the fixed conditioning approaches using weights $w\in\{0.6,0.8,1.0,1.2,1.4,2.0,2.5,3.0\}$. Results averaged across three random runs.
@@ 6.2 Baselines
We compare Diffusion-BBO with 10 baselines, including Bayesian optimization (BO), trust region BO (TuRBO) (Eriksson et al., 2019), local latent space Bayesian optimization (LOL-BO) (Maus et al., 2022), likelihood-free BO (LFBO) (Song et al., 2022), evolutionary algorithms (Brindle, 1980; Real et al., 2019), conditioning by adaptive sampling (CbAS) (Brookes et al., 2019), and random sampling. For BO approaches, we include Gaussian Processes (GPs) with Monte Carlo (MC)-based batch expected improvement (EI), MC-based batch upper confidence bound (UCB) (Wilson et al., 2017), and joint entropy search (JES (Hvarfner et al., 2022) as the acquisition functions. For LFBO, we use EI and probability of improvement (PI) as the acquisition functions.
@@ 6.3 Results
Figure 2 illustrates the performance across six datasets for all baselines and our proposed Diffusion-BBO. Notably, Diffusion-BBO utilizing UaE as the acquisition function consistently outperforms other baselines in both discrete and continuous settings, with the sole exception of the TF-BIND-8 task. These results align with our theoretical analysis, demonstrating the effectiveness of the designed UaE. Additionally, in the Ant and Dkitty tasks, Diffusion-BBO demonstrates a significant lead over all baseline methods, starting from the very first iteration of the online optimization process. This remarkable performance can be attributed to the conditional diffusion model of Diffusion-BBO, which can effectively learn the data manifold in the design space from the low-quality initial training dataset. In contrast, the forward approach employed by BO and LFBO, which relies solely on optimizing the forward surrogate model, is more prone to converging on suboptimal solutions.
@@ 6.4 Ablation study
In this section, we conduct ablation studies to investigate the impact of our designed acquisition function, UaE. We compare Diffusion-BBO with the fixed conditioning approach for data querying. Instead of using UaE to dynamically select $w$ and the corresponding $y_{k}=w\cdot\phi_{k}$ for conditional sampling, the fixed conditioning approach predetermines a fixed weight $w$ and consistently generates new samples conditioned on $w\cdot\phi_{k}$. As shown in Figure 3, Diffusion-BBO with UaE consistently outperforms the fixed conditioning approach. Furthermore, it can be found that simply conditioning on higher $w\cdot\phi_{k}$ by applying $w>1$ does not enhance optimization performance. This highlights the effectiveness of UaE in identifying the optimal $y_{k}$ for conditioning by balancing between targeting higher objective values and minimizing the corresponding epistemic uncertainty.
Additionally, we analyze the computational time for model training and acquisition function computation for Diffusion-BBO and existing baselines, as shown in Table 2. The results indicate that the computational time for Diffusion-BBO is comparable to BO approaches using GP as the surrogate model, typically ranging from 1 to 2 minutes per iteration for model training and acquisition function computation. Given the context of online BBO, where querying the oracle to generate new data is the most expensive or time-consuming part, a few minutes spent on training and acquisition function computation should not be considered a significant computational burden.
Furthermore, we evaluate the effect of batch size $N$ (i.e., the number of queries per iteration) on Diffusion-BBO. The results can be found in Appendix E.
Task GP (EI) GP (UCB) Evolution LFBO (EI) LFBO (PI) CbAS TuRBO Diffusion-BBO
TFBind8 $112.92$ $113.81$ $0.0021$ $0.59$ $1.44$ $0.075$ $67.23$ $136.29$
Molecular Discovery $53.14$ $53.82$ $0.0024$ $1.93$ $1.12$ $0.023$ $76.84$ $69.44$
Table 2: Model training and acquisition function computation time in seconds.
7 Discussion & Conclusion
@@
In this paper, we introduced Diffusion-BBO, a sample-efficient online BBO framework leveraging the conditional diffusion model as the inverse surrogate model. By utilizing the novel acquisition function UaE, Diffusion-BBO strategically proposes the conditioning variable $y$ in the objective space to improve the sample efficiency. Our empirical evaluations on $6$ real-world scientific discovery tasks demonstrate that Diffusion-BBO achieves state-of-the-art performance, establishing its potential as a robust tool for efficient online black-box optimization. Theoretically, we prove that using UaE leads to optimal optimization solutions. For future work, Diffusion-BBO can be extended to various Bayesian optimization settings, including multi-objective and multi-fidelity optimization, to develop an adaptive experimental design platform for scientific discovery tasks.
@@ Appendix B Analysis of Sub-optimality for Black-box Function
In this section, we study the behavior of the sub-optimality gap of our algorithm by proving Section 5.4 and Section 5.4. We first introduce the notation that is used throughout this section and the next section. Then we present the main lemmas along with their proofs. Finally, we combine the lemmas to prove our main results.
At each iteration $k\in[K]$, let $y_{k}^{*}$ be the target function value on which the diffusion model conditions, and $p_{\theta}$ be the model learned by the conditional diffusion model. We define the performance metric for online BBO problem, which measures the sub-optimal performance gap between the function value achieved by sample $\bm{x}\sim p_{\theta}(\cdot|y_{k}^{*},\mathcal{D})$ and the target function value $y_{k}^{*}$. Its formal definition is described as follows:
$\Delta(p_{\theta},y_{k}^{*})=\left|y_{k}^{*}-\max_{j\in[N]}f(\bm{x}_{j})\right|,~{}~{}\text{where}~{}~{}\bm{x}_{j}\sim p_{\theta}(\cdot|y_{k}^{*},\mathcal{D}),~{}~{}\forall j\in[N].$ (21)
For simplicity of analysis, we consider $N=1$, and let the generated sample at the $k$-th iteration be $\bm{x}_{k}$ in the remaining text. We remark that all proofs go through smoothly for general $N$ with more nuanced notations, and do not affect the conclusions being drawn. To proceed with the proofs in this section, we first state the formal assumptions for the black-box function $f(\cdot)$ and sample $\bm{x}$.
@@
Assumption 3.
Each generated sample $\bm{x}\in\mathbb{R}^{d}$ is $\sigma$-subGaussian. That is, there exists $\sigma\in\mathbb{R}$ such that for any $\bm{v}\in\mathbb{R}^{d}$ with $\left\|\bm{v}\right\|=1$, $\bm{v}^{\mathrm{T}}(\bm{x}-{\mathbb{E}}[\bm{x}])$ is $\sigma$-subGaussian, and its moment generating function is bounded by:
${\mathbb{E}}[\exp{\left(\lambda\bm{v}^{\mathrm{T}}(\bm{x}-{\mathbb{E}}[\bm{x}])\right)}]\leq\exp{\left(\frac{\sigma^{2}\lambda^{2}}{2}\right)},~{}~{}~{}~{}\forall\lambda\in\mathbb{R},~{}\bm{v}\in\mathbb{S}^{d-1},$
where $\mathbb{S}:=\{\bm{v}\in\mathbb{R}^{d}:\left\|\bm{v}\right\|=1\}$ is the $(d-1)$ unit sphere.
Before proceeding with the proofs of main theorems, we present our main lemmas.
Lemma B.1.
At each iteration $k\in[K]$, under fixed parameters $\theta$ and $\theta^{*}$, for $\bm{x}_{k}\sim p_{\theta}(\cdot|y_{k}^{*},\mathcal{D})$, $\bm{x}^{*}\sim p_{\theta^{*}}(\cdot|y_{k}^{*},\mathcal{D})$, we have
${\mathbb{E}}_{\bm{x}_{k}\sim p_{\theta}(\cdot|y_{k}^{*},\mathcal{D}),\bm{x}^{*}\sim p_{\theta^{*}}(\cdot|y_{k}^{*},\mathcal{D})}\left[\|\bm{x}^{*}-\bm{x}_{k}\|\right]$ $\leq 8\sqrt{d}\sigma+\left\|{\mathbb{E}}_{\bm{x}^{*}}[\bm{x}^{*}]-{\mathbb{E}}_{\bm{x}_{k}}[\bm{x}_{k}]\right\|,$ (22)
${\mathbb{E}}_{\bm{x}_{k}\sim p_{\theta}(\cdot|y_{k}^{*},\mathcal{D}),\bm{x}^{*}\sim p_{\theta^{*}}(\cdot|y_{k}^{*},\mathcal{D})}\left[\|\bm{x}^{*}-\bm{x}_{k}\|\right]$ $\geq\left\|{\mathbb{E}}_{\bm{x}^{*}}[\bm{x}^{*}]-{\mathbb{E}}_{\bm{x}_{k}}[\bm{x}_{k}]\right\|.$ (23)
Proof of Lemma B.1.
To bound ${\mathbb{E}}\left[\|\bm{x}^{*}-\bm{x}_{k}\|\right]$, by triangle inequality,
${\mathbb{E}}_{\bm{x}_{k},\bm{x}^{*}}\left[\|\bm{x}^{*}-\bm{x}_{k}\|\right]$ $={\mathbb{E}}\left[\|\bm{x}^{*}-{\mathbb{E}}[\bm{x}^{*}]+{\mathbb{E}}[\bm{x}_{k}]-\bm{x}_{k}+{\mathbb{E}}[\bm{x}^{*}]-{\mathbb{E}}[\bm{x}_{k}]\|\right]$
$\leq{\mathbb{E}}\left[\left\|\bm{x}^{*}-{\mathbb{E}}[\bm{x}^{*}]\right\|\right]+{\mathbb{E}}\left[\left\|\bm{x}_{k}-{\mathbb{E}}[\bm{x}_{k}]\right\|\right]+{\mathbb{E}}\left[\left\|{\mathbb{E}}[\bm{x}^{*}]-{\mathbb{E}}[\bm{x}_{k}]\right\|\right].$
Under 3, by Lemma B.3, we have,
${\mathbb{E}}_{\bm{x}_{k},\bm{x}^{*}}\left[\|\bm{x}^{*}-\bm{x}_{k}\|\right]\leq 8\sqrt{d}\sigma+\left\|{\mathbb{E}}[\bm{x}^{*}]-{\mathbb{E}}[\bm{x}_{k}]\right\|.$
Applying triangle inequality completes the step. In addition, it can be easily seen that
${\mathbb{E}}_{\bm{x}_{k},\bm{x}^{*}}\left[\|\bm{x}^{*}-\bm{x}_{k}\|\right]\geq\left\|{\mathbb{E}}[\bm{x}^{*}]-{\mathbb{E}}[\bm{x}_{k}]\right\|.$
Lemma B.2.
At each iteration $k\in[K]$, under fixed parameters $\theta$ and $\theta^{*}$, for $\bm{x}_{k}\sim p_{\theta}(\cdot|y_{k}^{*},\mathcal{D})$, $\bm{x}^{*}\sim p_{\theta^{*}}(\cdot|y_{k}^{*},\mathcal{D})$, we have
$\mathrm{Var}_{\bm{x}_{k}\sim p_{\theta}(\cdot|y_{k}^{*},\mathcal{D}),\bm{x}^{*}\sim p_{\theta^{*}}(\cdot|y_{k}^{*},\mathcal{D})}(\left\|\bm{x}^{*}-\bm{x}_{k}\right\|)\leq c_{3}d\sigma^{2}.$ (24)
Proof of Lemma B.2.
By definition of variance,
$\mathrm{Var}_{\bm{x}_{k},\bm{x}^{*}}(\left\|\bm{x}^{*}-\bm{x}_{k}\right\|)={\mathbb{E}}[\left\|\bm{x}^{*}-\bm{x}_{k}\right\|^{2}]-({\mathbb{E}}[\left\|\bm{x}^{*}-\bm{x}_{k}\right\|])^{2}.$ (25)
Expanding the first term leads to
${\mathbb{E}}[\left\|\bm{x}^{*}-\bm{x}_{k}\right\|^{2}]$ $={\mathbb{E}}[(\bm{x}^{*}-\bm{x}_{k})^{\mathrm{T}}(\bm{x}^{*}-\bm{x}_{k})]$
$={\mathbb{E}}[\left\|\bm{x}^{*}\right\|^{2}]+{\mathbb{E}}[\left\|\bm{x}_{k}\right\|^{2}]-2{\mathbb{E}}[(\bm{x}_{k})^{\mathrm{T}}\bm{x}^{*}]$
$={\mathbb{E}}[\left\|\bm{x}^{*}\right\|^{2}]+{\mathbb{E}}[\left\|\bm{x}_{k}\right\|^{2}]-2{\mathbb{E}}[(\bm{x}_{k})]^{\mathrm{T}}{\mathbb{E}}[\bm{x}^{*}],$ (26)
where the last equality is due to the independece between $\bm{x}^{*}$ and $\bm{x}_{k}$.
Under 3 and by Lemma B.4, we have
${\mathbb{E}}[\left\|\bm{x}^{*}\right\|^{2}]$ $={\mathbb{E}}[\left\|\bm{x}^{*}-{\mathbb{E}}[\bm{x}^{*}]+{\mathbb{E}}[\bm{x}^{*}]\right\|^{2}]$
$={\mathbb{E}}[(\bm{x}^{*}-{\mathbb{E}}[\bm{x}^{*}])^{\mathrm{T}}(\bm{x}^{*}-{\mathbb{E}}[\bm{x}^{*}])]+\left\|{\mathbb{E}}[\bm{x}^{*}]\right\|^{2}$
$=\mathrm{tr}({\mathbb{E}}[(\bm{x}^{*}-{\mathbb{E}}[\bm{x}^{*}])(\bm{x}^{*}-{\mathbb{E}}[\bm{x}^{*}])]^{\mathrm{T}})+\left\|{\mathbb{E}}[\bm{x}^{*}]\right\|^{2}$
$\leq Cd\sigma^{2}+\left\|{\mathbb{E}}[\bm{x}^{*}]\right\|^{2}.$
Here, the second equality holds as the cross terms vanish due to the fact that ${\mathbb{E}}[\bm{x}^{*}-{\mathbb{E}}[\bm{x}^{*}]]=0$. Similarly,
${\mathbb{E}}[\left\|\bm{x}_{k}\right\|^{2}]\leq Cd\sigma^{2}+\left\|{\mathbb{E}}[\bm{x}_{k}]\right\|^{2}.$
Substituting the above two results back to Equation 26,
${\mathbb{E}}[\left\|\bm{x}^{*}-\bm{x}_{k}\right\|^{2}]$ $\leq 2Cd\sigma^{2}+\left\|{\mathbb{E}}[\bm{x}_{k}]\right\|^{2}+\left\|{\mathbb{E}}[\bm{x}^{*}]\right\|^{2}-2{\mathbb{E}}[(\bm{x}_{k})^{\mathrm{T}}\bm{x}^{*}]$
$\leq 2Cd\sigma^{2}+\left\|{\mathbb{E}}[\bm{x}_{k}]-{\mathbb{E}}[\bm{x}^{*}]\right\|^{2}.$ (27)
Substituting Equation 27 back to Equation 25 and applying Lemma B.1 leads to
$\mathrm{Var}_{\bm{x}_{k},\bm{x}^{*}}(\left\|\bm{x}^{*}-\bm{x}_{k}\right\|)$ $\leq 2Cd\sigma^{2}+\left\|{\mathbb{E}}[\bm{x}_{k}]-{\mathbb{E}}[\bm{x}^{*}]\right\|^{2}-(8\sqrt{d}\sigma+\left\|{\mathbb{E}}[\bm{x}^{*}]-{\mathbb{E}}[\bm{x}_{k}]\right\|)^{2}\leq c_{3}d\sigma^{2}.$
With the above results, we are ready to prove Section 5.4 and Section 5.4.
@@
Recall that we consider the case where $N=1$, and denote $\bm{x}_{k}$ the generated sample in the $k$-th iteration, i.e. $\bm{x}_{k}\sim p_{\theta}(\cdot|y_{k}^{*},\mathcal{D})$, where $\theta\sim p(\theta\mid\mathcal{D})$. In each iteration $k$, with the existence of $\theta^{*}\sim p(\theta\mid\mathcal{D})$, we have $y_{k}^{*}=f(\bm{x}^{*})$, where $\bm{x}^{*}\sim p_{\theta^{*}}(\cdot|y_{k}^{*},\mathcal{D})$. Hence, under 2,
${\mathbb{E}}\left[{\Delta(p_{\theta},y_{k}^{*})}\right]={\mathbb{E}}\left[\left|f(\bm{x}^{*})-f(\bm{x}_{k})\right|\right]\leq L{\mathbb{E}}\left[\|\bm{x}^{*}-\bm{x}_{k}\|\right].$
By Lemma B.1, tower rule and Lemma B.5, we have
${\mathbb{E}}\left[{\Delta(p_{\theta},y_{k}^{*})}\right]$ $\leq L{\mathbb{E}}_{\theta,\theta^{*}}\left[{\mathbb{E}}_{x,x^{*}}\left[\|\bm{x}^{*}-\bm{x}_{k}\|\right]|\theta,\theta^{*}\right]$
$\leq 8L\sqrt{d}\sigma+{\mathbb{E}}_{\theta,\theta^{*}}\left[\left\|{\mathbb{E}}_{\bm{x}^{*}}[\bm{x}^{*}|\theta^{*}]-{\mathbb{E}}_{\bm{x}_{k}}[\bm{x}_{k}|\theta]\right\|\right]$
$\leq c_{1}L\sqrt{d}\sigma.$
@@
At every iteration $k\in[K]$, let the target function value on which the conditional diffusion model conditions be $y_{k}^{*}$. The statement needs to hold for each conditional diffusion model in the ensemble, and thus for simplicity of notation, the subscript $i$ of $\theta_{i}$ is dropped in the remaining proof. With the existence of $\theta^{*}\sim p(\theta\mid\mathcal{D})$, we have $y_{k}^{*}=f(\bm{x}^{*})$, where $\bm{x}^{*}\sim p_{\theta^{*}}(\cdot|y_{k}^{*},\mathcal{D})$. Recall that $f(\bm{x}_{k})$ is achieved by $\bm{x}_{k}\sim p_{\theta}(\cdot|y_{k}^{*},\mathcal{D})$, where $\theta\sim p(\theta\mid\mathcal{D})$, and $N=1$.
Thus, by Eve’s law, the overall variance of $\Delta(p_{\theta},y_{k}^{*})$ can be decomposed as:
$\mathrm{Var}\left(\Delta(p_{\theta},y_{k}^{*})\right)$ $=\mathrm{Var}\left(|y_{k}^{*}-f(\bm{x}_{k})|\right)$
$=\mathrm{Var}\left(|f(\bm{x}^{*})-f(\bm{x}_{k})|\right)$
$=\underbrace{{\mathbb{E}}_{\theta,\theta^{*}}\left[\mathrm{Var}_{\bm{x}_{k},\bm{x}^{*}}(|f(\bm{x}^{*})-f(\bm{x}_{k})|~{}\rvert~{}\theta,\theta^{*})\right]}_{T_{1}}+\underbrace{\mathrm{Var}_{\theta,\theta^{*}}({\mathbb{E}}_{\bm{x}_{k},\bm{x}^{*}}[|f(\bm{x}^{*})-f(\bm{x}_{k})|~{}\rvert~{}\theta,\theta^{*}])}_{T_{2}}.$
In particular, the first term $T_{1}$ corresponds to the aleatoric component and the second term $T_{2}$ corresponds to the episdemic component. We then proceed to bound the above two terms separately.
Step 1: bound $T_{1}$. Under 2,
$\mathrm{Var}_{\bm{x}_{k},\bm{x}^{*}}(|f(\bm{x}^{*})-f(\bm{x}_{k})|~{}\rvert~{}\theta,\theta^{*})\leq L^{2}\mathrm{Var}_{\bm{x}_{k},\bm{x}^{*}}(\left\|\bm{x}^{*}-\bm{x}_{k}\right\||~{}\rvert~{}\theta,\theta^{*}).$
Under 3 and by Lemma B.2,
$T_{1}\leq L^{2}{\mathbb{E}}_{\theta,\theta^{*}}[\mathrm{Var}_{\bm{x}_{k},\bm{x}^{*}}(\left\|\bm{x}^{*}-\bm{x}_{k}\right\||~{}\rvert~{}\theta,\theta^{*})]\leq c_{3}L^{2}d\sigma^{2}.$ (28)
Step 2: bound $T_{2}$. Under 2,
$T_{2}\leq L^{2}\mathrm{Var}_{\theta,\theta^{*}}({\mathbb{E}}_{\bm{x}_{k},\bm{x}^{*}}[\left\|\bm{x}^{*}-\bm{x}_{k}\right\||~{}\rvert~{}\theta,\theta^{*}]))$
By Lemma B.1,